OpenStudy (anonymous):

Solve: 16x2 − 81 = 0. x = 4, 9 x = −4, −9 x = 9 over 4, x = −9 over 4 x = 81 over 16, x = −81 over 16

Vocaloid (vocaloid):

hint: 16x^2 - 81 = A^2 - B^2 can you tell me what A and B are?

OpenStudy (abb0t):

@pooja195 can help!

OpenStudy (anonymous):

A and B are the numbers

Vocaloid (vocaloid):

right, but which numbers? 16x^2 - 81 = A^2 - B^2 what is A = ?

OpenStudy (anonymous):

- 9 over 4?

Vocaloid (vocaloid):

... what? 

Vocaloid (vocaloid):

so, if A^2 = 16x^2, what is A?

OpenStudy (anonymous):

A is 16x^2

Vocaloid (vocaloid):

nope, not quite... A^2 = 16x^2 A = ?

OpenStudy (anonymous):

A is 16^2

Vocaloid (vocaloid):

nope, try one more time... A^2 = 16x^2 A = square root of A^2 = ?

OpenStudy (anonymous):

ITS 100

OpenStudy (anonymous):

?

Vocaloid (vocaloid):

what is the square root of 16x^2?

OpenStudy (anonymous):

Well 256 is what 16 times 16 equals but its Square root would just be 16 ._.

Vocaloid (vocaloid):

not quite let's break it into two parts: 16 and x^2 what is the square root of 16?

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

right, now what's the square root of x^2?

OpenStudy (anonymous):

Thanks for the help I already firguered out the answer its B

OpenStudy (anonymous):

I mean C